OpenStudy (jokersmart):

Answer your answer in simplest form. (rewritten below) 8/a^2+1/a

OpenStudy (jokersmart):



OpenStudy (anonymous):

how would you add \[\frac{8}{25}+\frac{1}{5}\]?

OpenStudy (jokersmart):

That would be 13/25 I believe? c':

OpenStudy (anonymous):

yes, but the question wasn't really "what is the answer" but "how do you do it?"

OpenStudy (anonymous):

because you are going to do the same thing for \[\frac{8}{a^2}+\frac{1}{a}\] except with an \(a\) instead of a \(5\)

OpenStudy (jokersmart):

Whoops, I'm sorry about that ^^;

OpenStudy (anonymous):

i am sure you multiplied top and bottom of \(\frac{1}{5}\) by \(5\) to get \[\frac{5}{25}\] since that is the common denominator

OpenStudy (anonymous):

you have to do the same thing here \[\frac{8}{a^2}+\frac{1}{a^2}=\frac{8}{a^2}+\frac{a}{a^2}\]now add up top